OpenStudy (anonymous):

find the minimum y-value on the graph of y=f(x) f(x)=x^2+4x-1

OpenStudy (anonymous):

Put it in vertex form by completing the square.

OpenStudy (anonymous):

\[ x^2+bx = \left(x+\frac b 2\right)^2-\left(\frac b 2\right)^2 \] So \[ x^2+4x=(x+2)^2-4 \]And \[ y = ((x+2)^2-4)-1 = (x+2)^2-5\]

OpenStudy (anonymous):

This finally leaves us at \[ y-(-5)=(x-(-2))^2 \]So the vertex is at \((-2,-5)\).

OpenStudy (anonymous):



OpenStudy (anonymous):

The vertex must be the lowest point.

OpenStudy (anonymous):

@uniquejamaica get it?

OpenStudy (anonymous):

yes thank you